OpenStudy (anonymous):

Find the probability of rolling a sum of seven with a single toss of two fair six-sided dice.

OpenStudy (turingtest):

7=1+6=2+5=3+4 each die can take on one of these numbers so there are 6 possible configurations that make seven. The two dice have 6x6=36 total possible configurations, so 6/36=1/6 are your odds. (somebody check me)

OpenStudy (anonymous):

1/6 was right. thank you